OpenStudy (anonymous):

A firecracker shoots up from a hill 165 feet high with an initial speed of 110 feet per second. Using the formula H(t) = −16t2 + vt + s, determine how long it will take the firecracker to hit the ground. 7.2 seconds 7.5 seconds 7.8 seconds 8.1 seconds

OpenStudy (valpey):

You have an initial condition where H(0) = 165. This will help you solve for s. You also know that the initial velocity (v) = 110ft/s. You just need to solve for the final time (lets call it f) when H(f) = 0.

OpenStudy (valpey):

H(0) = -16(0)^2 + 110(0) + s = 165 Can you see what s must be? H(f) = -16(f)^2 + 110(f) + s = 0

OpenStudy (valpey):

I am assuming the firecracker lands somewhere 165 feet below the place it was fired from.

OpenStudy (anonymous):

I'm getting it so far.

OpenStudy (anonymous):

Wait so whats the answer?

OpenStudy (anonymous):

8.1 D

OpenStudy (anonymous):

Thank you!